Find the domain and range of the function (x,x/|x|).


1
Expert's answer
2022-05-23T13:38:06-0400
"f(x)=\dfrac{x}{|x|}"

"f(x)=\dfrac{x}{|x|}=\dfrac{x}{x}=1, x>0"

"f(x)=\dfrac{x}{|x|}=\dfrac{x}{-x}=-1, x<0"

"f(x)" is not defind at "x=0."


"Domain: (-\infin, 0)\cup (0, \infin)"


"Range: \{-1, 1\}"